OpenStudy (anonymous):

ODE y^2=cx+1/8 c^3 ------ y=2xy'+y^2(y')^3

OpenStudy (usukidoll):

??? \[y^2=cx+\frac{1}{8}c^3..............y=2xy'+y^2(y')^3\] or \[y^2=cx+\frac{1}{8c^3}..............y=2xy'+y^2(y')^3\]

OpenStudy (usukidoll):

I don't know what's going on.. can you go into more detail?

OpenStudy (anonymous):

\[y^2=cx+1/8 c^3 y=2xy'+y^2(y')^3\]

OpenStudy (anonymous):

y^2=cx+1/8 c^3 y=2xy'+y^2(y')^3 if the given left is a solution od the ODE right :))

OpenStudy (usukidoll):

so we are differentiating? one part of the equation with the fraction is still confusing.

OpenStudy (usukidoll):

so in the end we should have something like \[y=2xy'+y^2(y')^3 \]?

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

could you rewrite the first equation.. it's the fraction portion that is making me lost. I don't know if \[c^3y \] is in the denominator or the numerator

OpenStudy (anonymous):

its only c^3

OpenStudy (anonymous):

1/8 c^3

OpenStudy (usukidoll):

-_-! \[y^2=cx+\frac{1}{8}c^3. \]

OpenStudy (usukidoll):

omg I apologize the latex is crazy 

OpenStudy (usukidoll):

so we let the x and y be the variables and we treat c as a constant.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry